# Quantum automata and languages of finite index thanks:

### v1

@@ 6.1 The case of finite index matrix languages
Now we propose a first non trivial extension of the main results of decidability provided in [7]. Such an extension is based upon the machinery of Subsection 5, and some combinatorial arguments developed by Choffrut, Varricchio and the second author in [10] for the study of rational relations.
To this purpose, following [11], we first recall that a matrix context-free grammar (which we will henceforth simply call matrix grammar) is a tuple $G=\langle V,\Sigma,M,S\rangle$, where $V$, $\Sigma$, and $M$ are respectively the finite sets of nonterminals, terminals, and matrix rules, and $S\in V$ is the start symbol. Each matrix rule $m\in M$ is a finite sequence $m=(p_{1},\ldots,p_{s})$ where each $p_{i}$, $1\leq i\leq s$, is a context-free production from $V$ to $(V\cup\Sigma)^{*}$. For $m\in M$, we define $x\mathrel{\mathop{\Rightarrow}_{m}}y,x,y\in V^{*}$ if $x=x_{0}\mathrel{\mathop{\Rightarrow}_{p_{1}}}x_{1}\mathrel{\mathop{\Rightarrow}_{p_{2}}}\cdots\mathrel{\mathop{\Rightarrow}_{p_{s}}}x_{s}=y$, where $\mathrel{\mathop{\Rightarrow}_{p_{i}}}$ is the standard context-free derivation relation. Let $M^{*}$ be the set of the finite sequences of matrix rules and $\alpha\in M^{*}$. If $\alpha=\varepsilon$, then $x\Rightarrow_{\alpha}x,x\in V^{*}$; if one has
@@
with $x_{i}\in V^{*}$, $m_{j}\in M$, $1\leq j\leq s$, $0\leq i\leq s$, then we write $x_{0}\mathrel{\mathop{\Rightarrow}_{\alpha}}x_{s}$. The language generated by $G$, $L(G)=\{w\in\Sigma^{*}\mid S\mathrel{\mathop{\Rightarrow}_{\alpha}}w\}$. The following definition is instrumental.
@@
A matrix grammar $G=(V,\Sigma,M,S)$ is said to be restricted linear of index $k$, or simply restricted, if the set of nonterminals of $G$ has the form $V=\{S\}\cup V^{\prime}$, where $V^{\prime}=V_{1}\cup\dots\cup V_{k},$ and $V_{i}\cap V_{j}=\emptyset,$ for all $i\neq j$, and a matrix rule of $G$ can be only of one of the following forms
@@
Let $L$ be a language generated by a restricted grammar. Then one has:
@@
$\{\varphi(u_{1})\oplus\varphi(v_{1})^{\mbox{T}}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\mbox{T}}\mid(u_{1},v_{1},\ldots,u_{k},v_{k})\in C_{q}\}$ (30)
@@
$\ell(e)=\varphi(u_{1})\oplus\varphi(v_{1})^{\mbox{T}}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\mbox{T}}.$
@@
We now discuss some straightforward corollaries of Proposition 10. First, remark that, by applying the proposition to the matrix grammars of index k=1, we obtains [7], Prop. 22 (cf Corollary 2).
We recall that a bounded semilinear language is a finite union of linear languages which are languages of the form
@@
([7], Prop. 21) If $L$ is bounded semilinear then its closure ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Furthermore, if the quantum automaton ${\cal Q}$ is rational, the $(L,Q)$ Intersection problem is decidable.
@@
Let $L_{k}$, with $k\in\mathbb{N}$, be the language over $\Sigma=\{a,b\}$ given by $L_{k}=\{u^{k}:u\in\Sigma^{*}\}$. It is easily checked that $L_{k}$ can be generated by the restricted grammar $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$, and $M$ is given by the matrix rules $(S\rightarrow X_{1}\cdots X_{k})$, $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow\varepsilon)$, and, for every $\sigma\in\Sigma$, $(X_{1}\rightarrow\sigma X_{1},\ldots,X_{k}\rightarrow\sigma X_{k})$.
@@ 6.2 The case of monoidal languages
Let $G=\langle V,\Sigma,P,S\rangle$ be a monoidal grammar of index $1$. We rerun the proof of Proposition 9 for $G$. We proceed thus on $\mathop{\mathrm{Card}}\nolimits(V)$. If $\mathop{\mathrm{Card}}\nolimits(V)=1,$ i.e., $V=\{S\}$, then $G$ has a sole monoid of cycles $M_{S}$. Taking into account (26) and the form of the terminal productions of $G$, one has ${\mathbf{C}l}(\varphi(L_{S}))=\Psi({\mathbf{C}l}(M_{S}))$, where $\Psi(X\oplus Y):=XY^{\mbox{T}}.$ From this and the fact that, by hypothesis, ${\mathbf{C}l}(M_{S})$ is irreducible, it follows the claim in this case. The basis of induction is proved.
@@
$\Psi^{\prime\prime}(X\oplus Y\oplus Z):=X\Psi^{\prime}(Z)Y^{\mbox{T}}.$
@@
$\zeta(u,v)\ =\ (\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}})\ \oplus\ \varphi(L_{B_{1}})^{\mbox{T}}\cdots\varphi(L_{B_{t}})^{\mbox{T}})$ (35)
where, for every $i=1,\ldots,t$, $\varphi(L_{B_{i}})^{\mbox{T}}$ denotes the set of matrices
$\{m^{\mbox{T}}\ :\ m\in\varphi(L_{B_{i}})\}.$ (36)
@@
Indeed, let us check $M_{S}\subseteq\zeta(\widehat{C}_{S}).$ If $\varphi(u)\oplus\varphi(v)^{\mbox{T}}\in M_{S}$, then $S\displaystyle\mathop{\Rightarrow}^{*}uSv$. Since $G={\cal G}_{1}\circ{\cal G}_{2}$, the last derivation can be rewritten as
@@
$\varphi(u)\in\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}}),\quad\quad\varphi(v)^{\mbox{T}}\in\varphi(L_{B_{1}})^{\mbox{T}}\cdots\varphi(L_{B_{t}})^{\mbox{T}}.$ (39)
Finally observe that, by (37), $(\alpha,\beta)\in C_{S}$ so that $(\alpha,\beta^{\sim})\in\widehat{C}_{S}$. By (39) the latter implies $\varphi(u)\oplus\varphi(v)^{\mbox{T}}\in\zeta(\widehat{C}_{S})$. Hence $M_{S}\subseteq\zeta(\widehat{C}_{S})$. The other inclusion is checked similarly.
@@
${\mathbf{C}l}(\zeta({\cal R}))=\ ({\mathbf{C}l}(\varphi(L_{A_{1}}))\cdots{\mathbf{C}l}(\varphi(L_{A_{s}}))\ \oplus\ {\mathbf{C}l}(\varphi(L_{B_{1}}))^{\mbox{T}}\cdots{\mathbf{C}l}(\varphi(L_{B_{s}}))^{\mbox{T}}),$
taking into account that, for an arbitrary set $\cal A$ of orthogonal matrices, for the set (36), one has ${\mathbf{C}l}({\cal A}^{\mbox{T}})={\mathbf{C}l}({\cal A})^{\mbox{T}}$. By Lemma 8, the previous formula gives
${\mathbf{C}l}(\zeta({\cal R}))=\ (\Psi(H_{1})\cdots\Psi(H_{s})\ \oplus\ \Psi(G_{1})^{\mbox{T}}\cdots\Psi(G_{s})^{\mbox{T}}),$
@@
$\Psi^{\prime}(X_{1}\oplus\cdots\oplus X_{s}\oplus Y_{1}\oplus\cdots\oplus Y_{s})=(\Psi(X_{1})\cdots\Psi(X_{s})\oplus\Psi(Y_{1})^{\mbox{T}}\cdots\Psi(Y_{s})^{\mbox{T}}).$
@@ 7 Concluding remarks and open problems
[$\alpha$]  One could extend Proposition 11 to a broader class of quantum automata. In this context, the construction of the decidability procedure would require the computation of the Zariski closure of a group $G$ of matrices that is algebraically presented, i.e. $G$ admits a set of generators that is algebraic itself. It may be of interest to investigate conditions assuring the extension of the algorithms of [12, 17, 25] for such groups.
@@
[$\gamma$]  The $(L,Q)$ Intersection problem is decidable for bounded semi-linear languages. It may be of interest to get extensions of this result for the broader class of languages accepted by the reversal bounded non deterministic counter machines [18]. This extension could exploit a characterization of such languages provided by Ibarra in [19] in terms of some special combinatorial systems called RLGMC-grammars (right-linear grammar with multi-counters). Essentially, a RLGMC-grammar is a regular grammar that associates with each derivation a weight, i.e. a vector in the additive monoid $\mathbb{N}^{d}$. Then a derivation is considered successful if the corresponding weight belongs to the monoid generated by $\bf 1^{d}$ $=(1,1,\ldots,1)$. Here the main bulk is the construction of a formula (see Proposition 7) based on the computation of the Zariski closure of the Intersection of two finitely generated monoids of matrices. The intrinsic difficulty to control the accumulation points of the set – i.e., to exclude those one not coming from the effective functioning of the machines – makes such a computation a non trivial task.
@@ 8 Technical details and Examples
This implies that $\overline{\mbox{m}(\overline{{\cal A}},\overline{{\cal B}})}=\overline{\mbox{m}({\cal A},{\cal B})}.$ By using the very same argument for $H_{1}$ one verifies that $H_{2}$ is an irreducible effective algebraic set. Moreover, by (43) and (44), one has $H_{2}=\overline{\Psi(G)^{2}}$.

### v2

@@ 6.1 The case of finite index matrix languages
Now we propose a first non trivial extension of the main results of decidability provided in [7]. Such an extension is based upon the machinery of Section 5, and some combinatorial tools developed by Choffrut, Varricchio and the second author in [10] for the study of rational relations. To this purpose, following [11], Ch. $1$, Sec. $1.1$, we first recall that a matrix context-free grammar (which we will henceforth simply call matrix grammar) is a tuple $G=\langle V,\Sigma,M,S\rangle$, where $V$, $\Sigma$, and $M$ are respectively the finite sets of nonterminals, terminals, and matrix (of rules), and $S\in V$ is the start symbol. Each matrix $m\in M$ is a finite sequence $m=(p_{1},\ldots,p_{s})$ where each $p_{i}$, $1\leq i\leq s$, is a context-free production from $V$ to $(V\cup\Sigma)^{*}$. We denote by $P_{M}$ the set of all the context-free productions that appear in the matrices of $M$. For an arbitrary $m\in M$, we define the $1$-step derivation in $G$, $x\mathrel{\mathop{\Rightarrow}_{m}}y,$ with $x,y\in(V\cup\Sigma)^{*}$ if
$x=x_{0}\mathrel{\mathop{\Rightarrow}_{p_{1}}}x_{1}\mathrel{\mathop{\Rightarrow}_{p_{2}}}\cdots\mathrel{\mathop{\Rightarrow}_{p_{s}}}x_{s}=y,$
where $\mathrel{\mathop{\Rightarrow}_{p_{i}}}$ is the standard context-free derivation relation. This is equivalent to say that $x=x_{0},x_{s}=y$, and, for every $i=1,\ldots,s,$
$x_{i-1}=x^{\prime}_{i-1}X_{i}x^{\prime\prime}_{i-1},\quad x_{i}=x^{\prime}_{i-1}\alpha_{i}x^{\prime\prime}_{i-1},$
where $p_{i}=(X_{i}\rightarrow\alpha_{i})$, and for some $x^{\prime}_{i-1},x^{\prime\prime}_{i-1}\in(V\cup\Sigma)^{*}$. In other words, a $1$-step derivation in the matrix grammar $G$ corresponds to a $s$-step derivation in the context-free grammar $\langle V,\Sigma,P_{M},S\rangle$, i.e., using, one by one, all the context-free productions of $m$. Let $M^{*}$ be the set of the finite sequences of matrices and $\alpha\in M^{*}$. If $\alpha=\varepsilon$, then $x\Rightarrow_{\alpha}x,x\in V^{*}$; if one has
@@
with $x_{i}\in(V\cup\Sigma)^{*},$ $m_{j}\in M$, $1\leq j\leq s$, $0\leq i\leq s$, then we write $x_{0}\mathrel{\mathop{\Rightarrow}_{\alpha}}x_{s}$. The language generated by $G$ is $L(G)=\{w\in\Sigma^{*}\mid S\mathrel{\mathop{\Rightarrow}_{\alpha}}w\}$.
The following definition is instrumental.
@@
A matrix grammar $G=(V,\Sigma,M,S)$ is said to be restricted of index $k$, or simply restricted, if the set of nonterminals of $G$ has the form $V=\{S\}\cup V^{\prime}$, where $S\notin V^{\prime}$ and $V^{\prime}=V_{1}\cup\dots\cup V_{k},$ where $V_{i}\cap V_{j}=\emptyset,$ for all $i\neq j$, and a matrix of rules of $G$ can be only of one of the following forms
@@
Example 2
Let $L=\{a^{n}b^{n}c^{n}:n\in\mathbb{N}\}$ be the language over the alphabet $\Sigma=\{a,b,c\}$. Let $G$ be the restricted matrix grammar with the matrices of rules $m_{1}=(S\rightarrow ABC)$, $m_{2}=(A\rightarrow aA,B\rightarrow bB,C\rightarrow cC),$ and $m_{3}=(A\rightarrow~{}\varepsilon,B\rightarrow~{}\varepsilon,C\rightarrow\varepsilon).$ Note that the other components of $G$ can be deduced from these matrices. One easily checks that $L=L(G)$. Indeed, each word $a^{n}b^{n}c^{n},n\geq 0,$ can be generated by using first the matrix $m_{1}$, then the matrix $m_{2}$ $n-1$ times and finally the matrix $m_{3}$. Hence $L\subseteq L(G)$. The reverse inclusion is proved similarly.
@@
Let $L$ be a language generated by a restricted grammar and let $Q$ be a quantum automaton. Then one has:
@@
$\{\varphi(u_{1})\oplus\varphi(v_{1})^{\scriptstyle T}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\scriptstyle T}\mid(u_{1},v_{1},\ldots,u_{k},v_{k})\in C_{q}\}$ (30)
@@
$\ell(e)=\varphi(u_{1})\oplus\varphi(v_{1})^{\scriptstyle T}\oplus\cdots\oplus\varphi(u_{k})\ \oplus\ \varphi(v_{k})^{\scriptstyle T}.$
@@
We now discuss some straightforward corollaries of Proposition 10. First, remark that, by applying the proposition to the matrix grammars of index $k=1$, we obtains [7], Prop. 22 (cf Corollary 2). Moreover, the language of Example 2 is an instance of the following more general result. To this purpose, we recall that a bounded semilinear language is a finite union of linear languages which are languages of the form
@@
([7], Prop. 21) If $L$ is bounded semilinear then its closure ${\mathbf{C}l}(\varphi(L))$ is semialgebraic. Furthermore, if the quantum automaton ${\cal Q}$ is rational, then the $(L,Q)$ Intersection problem is decidable.
Example 2 (continued). Let ${\cal Q}=(s,\varphi,P,\lambda)$ be a finite quantum automaton where $\varphi:\Sigma^{*}\longrightarrow O_{n}$ is the morphism (4) associated with ${\cal Q}$. Recall that $\mathbf{C}l{({\cal A}))}$ and $\overline{{\cal A}}$ denote, respectively, the Euclidean and the Zariski closure of a set $\cal A$. The set ${\mathbf{C}l}(\varphi(L))$ is then equal to
${\mathbf{C}l}(\{\varphi(a)^{n}\varphi(b)^{n}\varphi(c)^{n}:n\in\mathbb{N}\})={\mathbf{C}l}(\Psi(\{\varphi(a)^{n}\oplus\varphi(b)^{n}\oplus\varphi(c)^{n}:n\in\mathbb{N}\}))$ (32)
where ()direct-sum($\oplus$)( ⊕ ) is the operation defined in (8) and $\Psi$ is the map defined, for every $X,Y,Z\in O_{n}$, as $\Psi(X\oplus Y\oplus Z):=XYZ$. By Theorem 1, the right-side term of (32) rewrites as $\Psi({\mathbf{C}l}(\{\varphi(a)^{n}\oplus\varphi(b)^{n}\oplus\varphi(c)^{n}:n\in\mathbb{N}\}))$, so that
${\mathbf{C}l}(\varphi(L))=\Psi({\mathbf{C}l}(\{g\}^{*}))=\Psi({\mathbf{C}l}(\langle\{g\}\rangle)),$
where $g=\varphi(a)\oplus\varphi(b)\oplus\varphi(c)$. By Theorem 3, ${\mathbf{C}l}(\langle\{g\}\rangle)=\overline{{\langle\{g\}\rangle}}$ is algebraic. From this, by Proposition 3, it follows that ${\mathbf{C}l}(\varphi(L))=\Psi(\overline{{\langle\{g\}\rangle}})$ is semialgebraic.
If the automaton $\cal Q$ is a rational, then, by applying Derksen’s algorithm to $\{g\}$, one computes $\overline{{\langle\{g\}\rangle}}$ as an algebraic set, which, in turn, implies that $\Psi(\overline{{\langle\{g\}\rangle}})$ is effective semialgebraic.
@@
Let $L_{k}$, with $k\in\mathbb{N}$, be the language over $\Sigma=\{a,b\}$ given by $L_{k}=\{u^{k}:u\in\Sigma^{*}\}$. It is easily checked that $L_{k}$ can be generated by the restricted grammar $G=\langle V,\Sigma,M,S\rangle$, where $V=\{S,X_{1},\ldots,X_{k}\}$, and $M$ is given by the matrix rules $(S\rightarrow X_{1}\cdots X_{k})$, $(X_{1}\rightarrow\varepsilon,\ldots,X_{k}\rightarrow~{}\varepsilon)$, and, for every $\sigma\in\Sigma$, $(X_{1}\rightarrow\sigma X_{1},\ldots,X_{k}\rightarrow\sigma X_{k})$. Note that the very same argument used in Example 2 allows one to verify that ${\mathbf{C}l}(\varphi(L))$ is semialgebraic and effective semialgebraic in the case that the quantum automaton is rational.
@@ 6.2 The case of monoidal languages
Let $G=\langle V,\Sigma,P,S\rangle$ be a monoidal grammar of index $1$. We rerun the proof of Proposition 9 for $G$. We proceed thus on $\mathop{\mathrm{Card}}\nolimits(V)$. If $\mathop{\mathrm{Card}}\nolimits(V)=1,$ i.e., $V=\{S\}$, then $G$ has a sole monoid of cycles $M_{S}$. Taking into account (26) and the form of the terminal productions of $G$, one has ${\mathbf{C}l}(\varphi(L_{S}))=\Psi({\mathbf{C}l}(M_{S}))$, where $\Psi(X\oplus Y):=XY^{\scriptstyle T}.$ From this and the fact that, by hypothesis, ${\mathbf{C}l}(M_{S})$ is irreducible, it follows the claim in this case. The basis of induction is proved.
@@
$\Psi^{\prime\prime}(X\oplus Y\oplus Z):=X\Psi^{\prime}(Z)Y^{\scriptstyle T}.$
@@
$\zeta(u,v)\ =\ (\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}})\ \oplus\ \varphi(L_{B_{1}})^{\scriptstyle T}\cdots\varphi(L_{B_{t}})^{\scriptstyle T})$ (36)
where, for every $i=1,\ldots,t$, $\varphi(L_{B_{i}})^{\scriptstyle T}$ denotes the set of matrices
$\{m^{\scriptstyle T}\ :\ m\in\varphi(L_{B_{i}})\}.$ (37)
@@
Indeed, let us check $M_{S}\subseteq\zeta(\widehat{C}_{S}).$ If $\varphi(u)\oplus\varphi(v)^{\scriptstyle T}\in M_{S}$, then $S\displaystyle\mathop{\Rightarrow}^{*}uSv$. Since $G={\cal G}_{1}\circ{\cal G}_{2}$, the last derivation can be rewritten as
@@
$\varphi(u)\in\varphi(L_{A_{1}})\cdots\varphi(L_{A_{s}}),\quad\quad\varphi(v)^{\scriptstyle T}\in\varphi(L_{B_{1}})^{\scriptstyle T}\cdots\varphi(L_{B_{t}})^{\scriptstyle T}.$ (40)
Finally observe that, by (38), $(\alpha,\beta)\in C_{S}$ so that $(\alpha,\beta^{\sim})\in\widehat{C}_{S}$. By (40) the latter implies $\varphi(u)\oplus\varphi(v)^{\scriptstyle T}\in\zeta(\widehat{C}_{S})$. Hence $M_{S}\subseteq\zeta(\widehat{C}_{S})$. The other inclusion is checked similarly.
@@
${\mathbf{C}l}(\zeta({\cal R}))=\ ({\mathbf{C}l}(\varphi(L_{A_{1}}))\cdots{\mathbf{C}l}(\varphi(L_{A_{s}}))\ \oplus\ {\mathbf{C}l}(\varphi(L_{B_{1}}))^{\scriptstyle T}\cdots{\mathbf{C}l}(\varphi(L_{B_{s}}))^{\scriptstyle T}),$
taking into account that, for an arbitrary set $\cal A$ of orthogonal matrices, for the set (37), one has ${\mathbf{C}l}({\cal A}^{\scriptstyle T})={\mathbf{C}l}({\cal A})^{\scriptstyle T}$. By Lemma 8, the previous formula gives
${\mathbf{C}l}(\zeta({\cal R}))=\ (\Psi(H_{1})\cdots\Psi(H_{s})\ \oplus\ \Psi(G_{1})^{\scriptstyle T}\cdots\Psi(G_{s})^{\scriptstyle T}),$
@@
$\Psi^{\prime}(X_{1}\oplus\cdots\oplus X_{s}\oplus Y_{1}\oplus\cdots\oplus Y_{s})=(\Psi(X_{1})\cdots\Psi(X_{s})\oplus\Psi(Y_{1})^{\scriptstyle T}\cdots\Psi(Y_{s})^{\scriptstyle T}).$
@@ 7 Concluding remarks and open problems
[$\alpha$]  One could extend Proposition 11 to a broader class of quantum automata. In this context, the construction of the decidability procedure would require the computation of the Zariski closure of a group $G$ of matrices that is algebraically presented, i.e. $G$ admits a set of generators that is algebraic itself. It may be of interest to investigate conditions ensuring the extension of the algorithms of [12, 17, 25] for such groups.
@@
[$\gamma$]  The $(L,Q)$ Intersection problem is decidable for bounded semi-linear languages. It may be of interest to get extensions of this result for the broader class of languages accepted by the reversal bounded non deterministic counter machines [18]. This extension could exploit a characterization of such languages provided by Ibarra in [19] in terms of some special combinatorial systems called RLGMC-grammars (right-linear grammar with multi-counters). Essentially, a RLGMC-grammar is a regular grammar that associates with each derivation a weight, i.e. a vector in the additive monoid $\mathbb{N}^{d}$. Then a derivation is considered successful if the corresponding weight belongs to the monoid generated by $\bf 1^{d}$ $=(1,1,\ldots,1)$. Here the main crux is the construction of a formula (see Proposition 7) based on the computation of the Zariski closure of the intersection of two finitely generated monoids of matrices. The intrinsic difficulty to control the accumulation points of the set – i.e., to exclude those one not coming from the effective functioning of the machines – makes such a computation a non trivial task.
@@ 8 Technical details and Examples
By using the very same argument for $H_{1}$ one verifies that $H_{2}$ is an irreducible effective algebraic set. Moreover, by (44) and (45), one has $H_{2}=\overline{\Psi(G)^{2}}$.